# Axionic quantum criticality of generalized Weyl semimetals

### v1

@@ Abstract
Signatures of the axionic insulation in Weyl materials have remained elusive so far, unless they are placed in strong magnetic fields [51]. Application of external magnetic fields quenches the kinetic energy and gives birth to Landau levels that are dispersive in the field direction. The resulting finite density of states triggers the onset of axion insulator even for infinitesimal interactions, following the magnetic catalysis mechanism [52, 53]. However, WSMs realized on strongly correlated quantum crystals, such as 227 pyrochlore iridates with multipolar magnetic orders [54, 55, 56, 57, 58] and Kondo materials [59, 60, 61], can be the ideal place to harness such exotic insulating ground state and hallmarks of emergent marginal Fermi liquids in terms of the logarithmic scaling corrections of physical observables, predicted in this work.
@@
Acknowledgments. This work was supported by ANID/ACT210100 (G.M., R.S.-G., and V.J.), the Swedish Research Council Grant No. VR 2019-04735 (V.J.), Fondecyt (Chile) Grants No. 1241033 (R.S.-G.) and No. 1230933 (V.J.). B.R. was supported by NSF CAREER Grant No. DMR-2238679 and is grateful to Institute for Solid State Physics, University of Tokyo where a part of this work was performed for their hospitality.

### v2

@@ Abstract
Unfortunately, signatures of the axionic insulation in Weyl materials have remained elusive so far, unless they are placed in strong magnetic fields [63]. Application of external magnetic fields quenches the kinetic energy and gives birth to Landau levels that are dispersive in the field direction. The resulting finite density of states triggers the onset of axion insulator even for infinitesimal interactions, following the magnetic catalysis mechanism [53, 64]. However, WSMs realized on strongly correlated quantum crystals, such as 227 pyrochlore iridates with multipolar magnetic orders [65, 66, 67, 68, 69] and Kondo materials [70, 71, 72], can be the ideal place to harness such exotic insulating ground state and hallmarks of emergent marginal Fermi liquids in terms of the logarithmic scaling corrections of physical observables, predicted in this work.
Acknowledgements.
@@
 This work was supported by ANID/ACT210100 (G.M., R.S.-G., and V.J.), the Swedish Research Council Grant No. VR 2019-04735 (V.J.), and Fondecyt (Chile) Grants No. 1241033 (R.S.-G.) and No. 1230933 (V.J.). B.R. was supported by NSF CAREER Grant No. DMR-2238679 and is grateful to the Institute for Solid State Physics, University of Tokyo where a part of this work was performed for their hospitality.
DATA AVAILABILITY No data were created or analyzed in this study.